OpenStudy (anonymous):

what is the solution to the given equation 2x + 4y = 12

Parth (parthkohli):

Many solutions.

Parth (parthkohli):

Diophantine equation?

OpenStudy (anonymous):

what is the solution of the equation? Given the linear equation

OpenStudy (anonymous):

2x + 4y =12

OpenStudy (goten77):

do you mean solve for y?

OpenStudy (goten77):

or x,y intercepts?

OpenStudy (anonymous):

x and y intercepts

Parth (parthkohli):

For the \(x\)-intercept, set \(y=0\) and solve for \(x\). For the \(y\)-intercept, set \(x=0\) and solve for \(y\).

OpenStudy (goten77):

hmm k then so x intercept is when y=0 2x+4y=12 becomes 2x+0=12 2x=12 <========just divided each side by 2 x=6 is the x intercept (6,0) y intercept is when x=0 2x+4y=12 becomes 0+4y=12 <=====just divided each side by 3 y=3 so the y intercept is at y=3 (0,3)

OpenStudy (anonymous):

thank you:)